OpenStudy (anonymous):

Find the greatest common factor of these two expressions. 27x^5 v^2 u^6 and 18w^4 v^7 I think the answer is 9^2 Any help? Please?

OpenStudy (pinksapphire):

But not every variable has a 9 or greater...

OpenStudy (anonymous):

oh oops I meant 3^2 ...3 cann go into 27 and 18

OpenStudy (whpalmer4):

\[27x^5v^2u^6\]\[18w^4v^7\] can't you factor out a 9 from both? How about \(v^2\)?

OpenStudy (pinksapphire):

But not every variable has 3 either...

OpenStudy (anonymous):

hmmm ok...can you teach me? I want to learn how to do this :)

OpenStudy (pinksapphire):

For example v^2 has a GCF of 1 because it's not 2v^2 or 3v^2

OpenStudy (anonymous):

Ok, that makes sense

OpenStudy (whpalmer4):

write each expression as its prime factors: \[27x^5v^2u^6 = 3*3*3*x*x*x*x*x*v*v*u*u*u*u*u*u = 3^3x^5v^2u^6\] \[18w^4v^7 = 2*3*3*w*w*w*w*v*v*v*v*v*v*v = 2^13^2w^4v^7\] the GCF is the highest common power of any factor found in both: GCF = \(3^2v^2 = 9v^2 \)

OpenStudy (whpalmer4):

You can factor \(3^2v^2\) out of both expressions: \[27x^5v^2u^6 = 9v^2(3x^5u^6)\]\[18w^4v^7 = 9v^2(2w^4v^5)\] there are no additional common factors which can be taken out after doing so, so \(9v^2\) is the GCF

OpenStudy (anonymous):

ohhh so I was close...just missing the v?

OpenStudy (whpalmer4):

yes, I wondered if that was a typo or a braino :-)

OpenStudy (anonymous):

Braino, I did half of the steps right...pffff always good to doubel check, I'm doing a review before I take the actual test.

OpenStudy (whpalmer4):

well, another way to think of this is to first find the GCF of the coefficients: GCF of 18 and 27 is obviously 9, right? Then find the GCF of the variables (should be pretty clear that is \(v^2\) here, I hope!) Multiply the GCF of variables and GCF of coefficients together to get the complete GCF

OpenStudy (anonymous):

ooooh thank you! all these neat tips and tricks :)